OpenStudy (anonymous):

Hi, please I need help with: if I have 2 lines with each one having x, y, z known coordinates, 1)how to find a third line orthogonal to both? 2)how to prove that there is a plane containing both lines, 3)if so what's the equation? thanks.

OpenStudy (amistre64):

a cross product

OpenStudy (amistre64):

there may not be a plane containing both lines unless the lines actually meet.

OpenStudy (anonymous):

Yes, I need to 1st prove that they intersect (the 2 lines) but...I don't know how to solve after that! how to find the 3rd line orthogonal to both...and the rest, please help!

OpenStudy (amistre64):

do you have any specific lines? or do youjust keep it general?

OpenStudy (anonymous):

yep...1st line: x=5-3t, y=2, z=2-t (parameter t) the 2nd line: x=s, y=3-4s, z=1+s (parameter s)

OpenStudy (amistre64):

x = 5 - 3t y = 2 + 0t z = 2 - 1t ^ ^ vector t = (-3,0,-1) point ----------------------------- x = 0 + 1s y = 3 - 4s z = 1 + 1s ^ ^vector s = (1, -4, 1) point does this make sense so far?

OpenStudy (anonymous):

yep, so far so good

OpenStudy (amistre64):

we can go a long route and try to compare point values, or we can go straight to defining a plane with at least one of the lines it it

OpenStudy (anonymous):

how to do that? could you please choose the simplest way to solve it?

OpenStudy (amistre64):

a cross product is a determinant of a matrix; in this case, a 3x3 matrix x y z x y -3, 0, -1 -3 0 1, -4, 1 1 -4 take the first 2 columns and repeat them; then run the diagonals - (0z 4x -3y) x y z x y -3, 0, -1 -3 0 1, -4, 1 1 -4 + (0x -1y 12z) add like terms -4x, 2y, 12z; our othogonal vectors is: (2,-1,-3) if i did that right

OpenStudy (amistre64):

(2,-1,-6) is better since i reduced it all by -2

OpenStudy (amistre64):

no we use that orthogonal vector and one of the point values to define a plane with one of the lines in it: lets use the top point (5,2,2) 2 (x-5) -1(y-2) -6(z-2) = 0

OpenStudy (anonymous):

oh! I never thought that algebra applies to this problem...

OpenStudy (amistre64):

it does :) my strategy was to define a plane that has the greatest chance of containing both lines, i know that one line is definantly in the plane, its only left to determine of the other xyz defnition fit into it, or if i only get a single solution for s

OpenStudy (amistre64):

using the xyz values of: x = 0 + 1s y = 3 - 4s z = 1 + 1s lets input them into the plane equation, and solve for s 2 (x-5) -1(y-2) -6(z-2) = 0 2 (0 + s - 5) -1(3 - 4s -2) -6(1+ s - 2) = 0 if s can only be one value, then this passes thru the plane and never is contained by it

OpenStudy (anonymous):

could you then find the equation???

OpenStudy (amistre64):

if the lines cross, then they will be defined for all values of t or s since they exist in one plane.  if they are parallel lines, the determinant we get would have been 0, or a normal vector of (0,0,0) which produces a true result regardless